# Global existence and boundedness in an attraction-repulsion chemotaxis system with nonlocal logistic source and sublinear productions

### v1

@@ 3 Global existence of classical solutions
On the other hand, using the Hölder’s inequality, we have
$\Big{\|}{u}\Big{\|}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq$ $C_{3}\Bigg{(}\int_{\Omega}{u}^{\theta p^{\prime}\frac{p+m-1}{\theta p^{\prime}}}\Bigg{)}^{\frac{\theta p^{\prime}}{p^{\prime}(p+m-1)}}\Bigg{(}\int_{\Omega}{u}^{(1-\theta)p^{\prime}\frac{1}{(1-\theta)p^{\prime}}}\Bigg{)}^{\frac{(1-\theta)p^{\prime}}{p^{\prime}}}$
$\leq$ $C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{(1-\theta)}_{{\mathcal{L}^{1}}(\Omega)},$
where $\theta=\frac{(p^{\prime}-1)(p+m-1)}{p^{\prime}(p+m-2)}\in(0,1)$. Similar way, we pick up
$\Big{\|}{u}\Big{\|}_{{\mathcal{L}^{p}}(\Omega)}\leq$ $C_{4}\Bigg{(}\int_{\Omega}{u}^{\eta p\frac{p+m-1}{\eta p}}\Bigg{)}^{\frac{\eta p}{p(p+m-1)}}\Bigg{(}\int_{\Omega}{u}^{(1-\eta)p\frac{1}{(1-\eta)p}}\Bigg{)}^{\frac{(1-\eta)p}{p}}$
$\leq$ $C_{4}\Big{\|}{u}\Big{\|}^{\eta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{(1-\eta)}_{{\mathcal{L}^{1}}(\Omega)},$
with $\eta=\frac{(p-1)(p+m-1)}{p(p+m-2)}\in(0,1)$. Hence, we obtain the following estimates from the above two inequalities,
$(i)\quad C_{2}\Big{\|}{u}\Big{\|}^{\frac{pa}{2}}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq C_{2}\Bigg{(}C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\theta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{\frac{pa}{2}}\leq C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)},$ (3.15)
$(ii)\quad C_{1}\Big{\|}{u}\Big{\|}^{\gamma(p+m-1)}_{{\mathcal{L}^{p}}(\Omega)}\leq C_{1}\Bigg{(}C_{4}\Big{\|}{u}\Big{\|}^{\eta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\eta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{\gamma(p+m-1)}\leq C_{6}\Big{\|}{u}\Big{\|}^{\eta\gamma(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)},$ (3.16)
$(iii)\quad\Big{\|}{u}\Big{\|}^{(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq\Bigg{(}C_{3}\Big{\|}{u}\Big{\|}^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big{\|}{u}\Big{\|}^{1-\theta}_{{\mathcal{L}^{1}}(\Omega)}\Bigg{)}^{(1-\gamma)(p+m-1)}\leq C_{7}\Big{\|}{u}\Big{\|}^{\theta(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}.$ (3.17)
Substituting (3.15)-(3.17) in (3.14), finally we achieve
@@
$\leq$ $\frac{2d_{1}(p-1)}{p^{2}\mu}\Big{\|}{\nabla{u}}^{\frac{p}{2}}\Big{\|}^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{5}\Big{\|}{u}\Big{\|}^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{8}\Big{\|}{u}\Big{\|}^{[\eta\gamma+\theta(1-\gamma)](p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}.$ (3.18)
Again, using the Young’s inequality, the second term can be written as
@@
By simple calculations, we guarantee that $\frac{pa\theta}{2(p+m-1)}<1$. That is,
$\frac{(1-\gamma)(p+m-1)}{1-\frac{\gamma(p+m-1)}{p}}\frac{1-\frac{1}{p^{\prime}}}{1-\frac{1}{p+m-1}}\frac{1}{p+m-1}<$ 11 11
$(1-\gamma)(p+m-1)\left(1-\frac{1}{p^{\prime}}\right)<$ $\left(1-\frac{\gamma(p+m-1)}{p}\right)(p+m-2)$
$\frac{1}{p+m-1}+\frac{\gamma-1}{p^{\prime}}<$ $\gamma\left(1-\frac{p+m-1}{p}\right)+\frac{\gamma}{p}$
$\frac{p}{\gamma(p+m-1)}-\frac{p}{\gamma p^{\prime}}<$ $(2-m)-\frac{p}{p^{\prime}}$
@@
Once again, using the Young’s inequality to the third term of (3.18), the estimates gives
@@
Now (3.19) holds true for the case $n=2$ whenever $m<2-\frac{2}{r}<2$ and also if, $m<2$, (3.19) remains true for the case $n=1$. This completes the proof. ∎
@@
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 1$ is an open, bounded domain with smooth boundary. Moreover, let $f$ fulfills (1.10) and for any $p>1$ and $M_{0}>0$, if $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$, then we have
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{w}}+\mu p\int_{\Omega}{{u}}^{p+m-1}-\mu pM_{0}\int_{\Omega}{{u}}^{p+m-1}.$ (3.24)
@@
where $C_{1}=\left(\epsilon\delta\right)^{\frac{-1}{\delta-1}}\frac{\delta-1}{\delta}\left(\chi p\frac{p-1}{p+k-1}\right)^{\frac{\delta}{\delta-1}}$. Similarly for $I_{2}$
@@
where $C_{2}=\left(\epsilon\delta\right)^{\frac{-1}{\delta-1}}\frac{\delta-1}{\delta}\left(\xi p\frac{p-1}{p+k-1}\right)^{\frac{\delta}{\delta-1}}$. Substituting. (3.26) and (3.27) in (3.25), we obtain
@@
$\int_{\Omega}{{u}}^{p}$ $=\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{2}}(\Omega)}\leq C_{3}\left(\left\|\nabla{u}^{\frac{p}{2}}\right\|^{2b}_{{\mathcal{L}^{2}}(\Omega)}\,\,\left\|{u}^{\frac{p}{2}}\right\|^{2(1-b)}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}+\left\|{u}^{\frac{p}{2}}\right\|^{2}_{{\mathcal{L}^{\frac{2}{p}}}(\Omega)}\right)$
@@
where $b=\frac{\frac{p}{2}-\frac{1}{2}}{\frac{p}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$. Now, we can estimate the second term in (3.28) using the above inequality as follows
$\frac{np+2}{2-(k-1)n}\int_{\Omega}{{u}}^{p}$ $\leq\frac{d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{6}.$ (3.29)
Again, using the Gagliardo-Nirenberg inequity, to get
@@
$\leq$ $C_{8}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{9}.$
Now, we can estimate the third term in (3.28) using the above inequality as follows
$\frac{-d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}\leq\frac{-d_{1}(p-1)}{pC_{8}}\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{10}.$ (3.30)
Substituting (3.29) and (3.30) in (3.28), we arrive at
@@
$+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}+C_{11}.$ (3.31)
Applying the variation-of-constants formula to (3.31), we obtain the following estimate with $\epsilon=\frac{d_{1}(p-1)}{4pC_{8}}$
@@
where $C_{12}=\int_{\Omega}{u}(\cdot,s_{0})+C_{11}\frac{2-(k-1)n}{np+2}$. From the Lemma 2.3, there exists $C_{p}>0$ such that
@@
where $C_{18}=C_{12}+C_{14}\Big{\|}{v}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}+C_{16}\Big{\|}{w}(\cdot,s_{0})\Big{\|}^{\frac{np+2}{2-(k-1)n}}_{{\mathcal{W}}^{2,\frac{np+2}{2-(k-1)n}}}$. Using the Young’s inequality with $l+k<1+\frac{2}{n}$, we have
@@
where $C_{19}=\left(\frac{d_{1}(p-1)}{2pC_{8}}\frac{\frac{2}{n}+1-k}{l}\right)^{\frac{-l}{\frac{2}{n}+1-(k+l)}}\frac{\frac{2}{n}+1-(k+l)}{\frac{2}{n}+1-k}C_{17}^{\frac{\frac{2}{n}+1-k}{\frac{2}{n}+1-(k+l)}}|\Omega|$. Substituting (3.36) in (3.35), finally, we arrive at
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}{\nabla{w}}.$
@@
Proof of Theorem 1.1. We can argue the proof by contradiction. Suppose that $T_{\mathrm{max}}<\infty$, utilizing the Lemma 2.4, the bounds (2.9), (2.10) and (2.16) holds for $t\in(s_{0},T_{\mathrm{max}})$. The blow-up criterion (2.1) contradicts with the above result that is, the maximum time is not finite. Finally, we done the following bounds with (3.4)
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq C_{17},\qquad\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\leq C_{18}\qquad\mathrm{and}\qquad\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\leq C_{19},$
for all $t>0$. This completes the proof.
4 Numerical Examples
This section discusses the results of the numerical simulations that clarify the behaviour of the solutions of the attraction-repulsion chemotaxis system in two space dimensions based on the finite difference method. In all of the examples below, we consider $d_{1}=d_{2}=d_{3}=\chi=\xi=\alpha=\beta=1$ and use a uniform grid $\Delta x=\Delta y=\frac{1}{201}$ in a square domain $[-0.5,0.5]\times[-0.5,0.5]$. Moreover, in the following examples, $\|u-u^{*}\|_{\infty}$ represents the maximum of difference between present time and previous time of $u(x,y,t)$.
Example 4.1 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})}\qquad\mbox{and}\qquad v(x,y,0)=w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
The attraction-repulsion chemotaxis system is tested in this example with and without a nonlocal source under same initial conditions for both $v$ and $w$. First we compute the cell densities for $\mu=0$ at a large time $t=10$ which are plotted in Figure 1.
((a))
((b))
((c))
Figure 1: Cell and chemical concentrations for $\mu=0$ at $t=10$.
Now compute the cell densities for $\mu=1$ for a large time $t=10$ and the obtained cell densities plotted in Figure 2.
((a))
((b))
((c))
Figure 2: Cell and chemical concentrations for $\mu=1$ at $t=10$.
From the above two situations, we observe that the solution decays in time and behaves like the fundamental solution of a heat equation. This is because the repulsion nullifies the attraction due to same initial conditions.
From the above two tests, we conclude that the nonlocal source term substantially reduces the convergence range of the solution and this source term may help to delay the blowup time. Moreover from the above discussions, we notice that when both the chemical concentrations $v$ and $w$ are initially the same, the solution cannot blowup. This numerical example particularly confirms the theoretical results of Jin[21].
Example 4.2 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=500e^{-50(x^{2}+y^{2})},\quad w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
Similar to the previous case, one can observe the convergence here. Hence it is concluded that whenever the attraction dominates the repulsion with an appropriate initial data, the solution converges for both the cases of $\mu=0$ and $1$.
((a))
((b))
((c))
Figure 3: Cell and chemical concentrations for $\mu=0$ at $t=10$.
((d))
((e))
((f))
Figure 4: Cell and chemical concentrations for $\mu=1$ at $t=10$.
Example 4.3 (Quick blowup of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=250e^{-25(x^{2}+y^{2})},\quad w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
This example deals with the behavior of the solution when the concentration of $w$ is initially higher than that of $v$. Figure 6 represents the cell densities at $t=3\times 10^{-5}$ for $\mu=0$. Unfortunately the solution immediately blows up after the time $t=2.5\times 10^{-5}$ and the spiky shape of the cell concentration is collapsed. Additionally we test the same system with a nonlocal source, the captured cell densities are figured out at time $t=3\times 10^{-5}$ in Figure 6.
((a))
((b))
Figure 5: Behaviour of $u$ for $\mu=0$ at $t=3\times 10^{-5}$.
((c))
((d))
Figure 6: Behaviour of $u$ for $\mu=1$ at $t=3\times 10^{-5}$.
From the two cases, one can see that the cells proliferate uncontrollably in the domain since the repulsion dominates the attraction.
Example 4.4 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u$, $k=m=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=100e^{-10(x^{2}+y^{2})},\quad v(x,y,0)=25e^{-2.5(x^{2}+y^{2})},\quad w(x,y,0)=50e^{-5(x^{2}+y^{2})}.$
This example deals with the behaviour of the system (1.6) when repulsion dominates the attraction under a small initial data. Contrary to the previous Example 4.3, the solution converges with an appropriate small initial data. From this, we conclude that whenever repulsion dominates the attraction, the small initial data help to attain the convergence; this can be seen in Figure 8.
((a))
((b))
((c))
((a))
((b))
((c))
Figure 8: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Example 4.5 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1,m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})}\qquad\mbox{and}\qquad v(x,y,0)=w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
Compared to solutions in Example 4.1, here the solution rather converges in a very short time and the exponent $m$ plays a crucial role in this convergence. This can be seen in the following Figure 9.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 9: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Example 4.6 (Convergence of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$, $\mu=1$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=500e^{-50(x^{2}+y^{2})},\quad w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
Figure 13 shows the cell densities at $t=10^{-4}$. One can see from the solution at time $t=10^{-4}$ that the cell density has almost reached the blowup state. But fortunately at time $t=10^{-3}$, the solution start to converge and the maximum value of $u$ keeps decreasing compared to the previous time $t=10^{-4}$. Now while we go for the next time $t=10^{-2}$, the solution converges absolutely and the growth of the cell density reaches equilibrium.
((a))
((b))
((c))
Figure 10: Behaviour of $u$ at $t=10^{-4}$.
((d))
((e))
((f))
Figure 11: Behaviour of $u$ at $t=10^{-3}$.
((a))
((b))
((c))
Figure 12: Behaviour of $u$ at $t=10^{-2}$.
((a))
((b))
((c))
Figure 13: Cell and chemical concentrations at $t=10$.
This experiment shows that, for specific values of the exponents $k,l$ and $m$, even if the initial condition is large, the system can be prevented from the blowup.
Example 4.7 (Quick blowup of solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\quad v(x,y,0)=250e^{-25(x^{2}+y^{2})},\quad w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
Quite similar to Example 4.3, the following figures indicate the cell densities captured at $t=3\times 10^{-5}$. Note that the solution seems to converge but at much later time $t=3\times 10^{-5}$, the solution suddenly blows up and the spiky shape is collapsed.
((a))
((b))
Figure 14: Behaviour of $u$ for $\mu=0$ at $t=3\times 10^{-5}$.
((c))
((d))
Figure 15: Behaviour of $u$ for $\mu=1$ at $t=3\times 10^{-5}$.
From this example, it is observed that the solution still blows up for a higher value of $m$ due to the domination of attraction by repulsion.
Example 4.8 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.5$, $k=1$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=100e^{-10(x^{2}+y^{2})},\quad v(x,y,0)=25e^{-2.5(x^{2}+y^{2})}\quad\mbox{and}\quad w(x,y,0)=50e^{-5(x^{2}+y^{2})}.$
This example deals with the behaviour of the system (1.6) when repulsion dominates the attraction for particular values of the exponents $k,l$ and $m$ under a small initial data. Unlike the previous Example 4.7, the solution converges with a small initial data.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 16: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
The system (1.6) admits imaginary solutions for values of $k$ greater than 1 and the captured cell densities are depicted in the examples below for various scenarios.
Example 4.9 (Imaginary solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\>v(x,y,0)=500e^{-50(x^{2}+y^{2})},\>w(x,y,0)=250e^{-25(x^{2}+y^{2})}.$
The behavior of the solution when $v>w$ initially, is illustrated in the figures below. It can be observed that for $\mu=0$, the solution transits to an imaginary state after $t=0.76\times 10^{-6}$, while for $\mu=1$, this transition occurs after $t=0.77\times 10^{-6}$.
((a))
((b))
Figure 17: Behaviour of $u$ for $\mu=0$ at $t=0.76\times 10^{-6}$.
((c))
((d))
Figure 18: Behaviour of $u$ for $\mu=1$ at $t=0.77\times 10^{-6}$.
Example 4.10 (Imaginary solution).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=1000e^{-100(x^{2}+y^{2})},\>v(x,y,0)=250e^{-25(x^{2}+y^{2})},\>w(x,y,0)=500e^{-50(x^{2}+y^{2})}.$
The subsequent figures illustrate how the solution behaves when initial values $v<w$. In this scenario, it is observed that for $\mu=0$, the solution reaches an imaginary state after $t=0.74\times 10^{-6}$ and for $\mu=1$, this transitions occurs after $t=0.75\times 10^{-6}$.
((a))
((b))
Figure 19: Behaviour of $u$ for $\mu=0$ at $t=0.74\times 10^{-6}$.
((c))
((d))
Figure 20: Behaviour of $u$ for $\mu=1$ at $t=0.75\times 10^{-6}$.
Example 4.11 (Convergence of solution for small initial data).
Consider the attraction-repulsion chemotaxis system (1.6) with $f(u)=u^{l}$, where $l=0.4$, $k=1.5$, $m=1.8$ subject to the radially symmetric bell-shaped initial data
$u(x,y,0)=10e^{-10(x^{2}+y^{2})},\>v(x,y,0)=5e^{-5(x^{2}+y^{2})},\>w(x,y,0)=2.5e^{-2.5(x^{2}+y^{2})}.$
The behavior of the solution is illustrated in the following figures, assuming the initial values $v>w$. It is observed that with small initial data, the solution converges for $k=1.5$. These figures demonstrate the long-term behavior of the system (1.6) at $t=10$ for $\mu=0$ and $\mu=1$.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 21: Cell and chemical concentrations for $\mu=0$ (first row) and $\mu=1$ (second row) at $t=10$.
Remark 4.1.
If $k>1$, using our approach as in the above Example 4.11, the convergence can be achieved for the cases $v(x,y,0)=w(x,y,0)$ and $v(x,y,0)<w(x,y,0)$.

### v2

@@ 3 Global existence of classical solutions
On the other hand, using the Hölder’s inequality with $1<p^{\prime}<p+m-1$, we have
$\Big\|{u}\Big\|_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq$ $C_{3}\Bigg(\int_{\Omega}{u}^{\theta p^{\prime}\frac{p+m-1}{\theta p^{\prime}}}\Bigg)^{\frac{\theta p^{\prime}}{p^{\prime}(p+m-1)}}\Bigg(\int_{\Omega}{u}^{(1-\theta)p^{\prime}\frac{1}{(1-\theta)p^{\prime}}}\Bigg)^{\frac{(1-\theta)p^{\prime}}{p^{\prime}}}\leq C_{3}\Big\|{u}\Big\|^{\theta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big\|{u}\Big\|^{(1-\theta)}_{{\mathcal{L}^{1}}(\Omega)},$
where $\theta=\frac{(p^{\prime}-1)(p+m-1)}{p^{\prime}(p+m-2)}\in(0,1)$ and $C_{3}>0$. Hence, we obtain
$C_{2}\Big\|{u}\Big\|^{\frac{pa}{2}}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq C_{4}\Big\|{u}\Big\|^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}\qquad\text{and}\qquad\Big\|{u}\Big\|^{(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p^{{}^{\prime}}}}(\Omega)}\leq C_{5}\Big\|{u}\Big\|^{\theta(1-\gamma)(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}.$ (3.14)
Similar way, we obtain
$\Big\|{u}\Big\|_{{\mathcal{L}^{p}}(\Omega)}\leq$ $C_{6}\Bigg(\int_{\Omega}{u}^{\eta p\frac{p+m-1}{\eta p}}\Bigg)^{\frac{\eta p}{p(p+m-1)}}\Bigg(\int_{\Omega}{u}^{(1-\eta)p\frac{1}{(1-\eta)p}}\Bigg)^{\frac{(1-\eta)p}{p}}\leq C_{6}\Big\|{u}\Big\|^{\eta}_{{\mathcal{L}^{p+m-1}}(\Omega)}\Big\|{u}\Big\|^{(1-\eta)}_{{\mathcal{L}^{1}}(\Omega)},$
$\leq$ $C_{7}\Big\|{u}\Big\|^{\eta\gamma(p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)}$ (3.15)
with $\eta=\frac{(p-1)(p+m-1)}{p(p+m-2)}\in(0,1)$ and $C_{i}>0,i=4,5,6,7$. Substituting (3.14) and (3.15) in (3.13), finally we achieve
@@
$=$ $\frac{2d_{1}(p-1)}{p^{2}\mu}\Big\|{\nabla{u}}^{\frac{p}{2}}\Big\|^{2}_{{\mathcal{L}^{2}}(\Omega)}+C_{5}\Big\|{u}\Big\|^{\frac{pa\theta}{2}}_{{\mathcal{L}^{p+m-1}}(\Omega)}+C_{8}\Big\|{u}\Big\|^{[\eta\gamma+\theta(1-\gamma)](p+m-1)}_{{\mathcal{L}^{p+m-1}}(\Omega)},$ (3.16)
where $C_{8}>0$. Again, using the Young’s inequality with $C_{9},C_{10}>0$, the second term can be written as
@@
By simple computations, we calculate
@@
This guarantee that $\frac{pa\theta}{2(p+m-1)}<1$. Once again, using the Young’s inequality with $C_{11},C_{12}>0$ to the third term of (3.16), the estimates gives
@@
Now (3.18) holds true for the case $n=2$ whenever $m<2-\frac{2}{r}$. This completes the proof. ∎
@@
Suppose that $\Omega\subset\mathbb{R}^{n}\>(n\geq 2)$, is a bounded domain with smooth boundary. Moreover, let $f$ fulfills (1.10) and for any $p>1$, if $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$, then we have
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}\cdot{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}\cdot{\nabla{w}}+\mu p\int_{\Omega}{{u}}^{p+m-1}-\mu pM_{0}\int_{\Omega}{{u}}^{p+m-1}.$ (3.22)
@@
where $C_{1}>0$. Similarly for $I_{2}$
@@
where $C_{2}>0$. Substituting. (3.24) and (3.25) in (3.23), we obtain
@@
[…]
@@
where $b=\frac{\frac{p}{2}-\frac{1}{2}}{\frac{p}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$ and $C_{5}>0$. Now, we can estimate the second term in (3.26) using the above inequality as follows
$\frac{np+2}{2-(k-1)n}\int_{\Omega}{{u}}^{p}$ $\leq\frac{d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{6},$ (3.27)
where $C_{6}>0$. Again, using the Gagliardo-Nirenberg inequity, to get
@@
$\leq$ $C_{8}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}+C_{9},$
where $C_{8},C_{9}>0$ Now, we can estimate the third term in (3.26) using the above inequality as follows
$\frac{-d_{1}(p-1)}{p}\int_{\Omega}\left|\nabla{u}^{\frac{p}{2}}\right|^{2}\leq\frac{-d_{1}(p-1)}{pC_{8}}\int_{\Omega}{u}^{\frac{np+2}{n}}+C_{10},$ (3.28)
where $C_{10}>0$. Substituting (3.27) and (3.28) in (3.26), we arrive at
@@
$+C_{1}\int_{\Omega}{|\Delta{v}|}^{\frac{np+2}{2-(k-1)n}}+C_{2}\int_{\Omega}{|\Delta{w}|}^{\frac{np+2}{2-(k-1)n}}+C_{11},$ (3.29)
where $C_{11}>0$. Applying the variation-of-constants formula to (3.29), we obtain the following estimate with $\epsilon=\frac{d_{1}(p-1)}{4pC_{8}}$
@@
where $C_{12}>0$. From Lemma 2.3, there exists $C_{p}>0$ such that
@@
where $C_{18}>0$. Using the Young’s inequality with $l+k<1+\frac{2}{n}$, we have
@@
where $C_{19}>0$. Substituting (3.34) in (3.33), finally, we arrive at
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{{u}}^{p}\leq$ $-d_{1}p(p-1)\int_{\Omega}{{u}}^{p-2}|{\nabla{u}}|^{2}+\chi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}\cdot{\nabla{v}}$
$-\xi p(p-1)\int_{\Omega}{u}^{p+k-2}{\nabla{u}}\cdot{\nabla{w}}.$
@@
Proof of Theorem 1.1. We proceed by contradiction. Assume that $T_{\mathrm{max}}<\infty$. By Lemma 2.4, the bounds remain valid for all $t\in(0,T_{\mathrm{max}})$. However, the blow-up criterion (2.1) contradicts this result. Hence, the maximal existence time cannot be finite i.e, $T_{\mathrm{max}}=\infty$. Consequently, we obtain the uniform estimates
$\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+\big\|{v}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}+\big\|{w}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}\leq C,$
for all $t>0$ and the constant $C$ is positive. This completes the proof.